The sum of the digits of a certain two digits number is 6. If the original number is subtracted from the number formed by interchanging the digits, the results is 36. Find the original number. Please use U for units digit and T for tens digit
Answer by richwmiller(17219)

u+t=6,
(10u+t)-(10t+u)=36
9u-9t=36
u-t=4
2u=10
u=5
t = 1, u = 5